OpenStudy (anonymous):

Enter the value that completes this sentence:The distance between points A and D is _____ units.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

the d isnt part of anything

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

do you know how to do it?

OpenStudy (aripotta):

by just looking at the graph, you can see the distance between a and d

OpenStudy (aripotta):

however, you can also use the distance formula if you want:\[\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\]

OpenStudy (anonymous):

dang sorry i copied the wrong question

OpenStudy (anonymous):

ok here let me get the right question up

OpenStudy (aripotta):

ok, use the distance formula \[\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\]

OpenStudy (anonymous):

ok

OpenStudy (aripotta):

what do you get?

OpenStudy (anonymous):

\[\sqrt{(2-(-4)^{2}+(-4 -1)^{2}}\]

OpenStudy (anonymous):

like that

OpenStudy (aripotta):

yes. keep working to get the answer

OpenStudy (anonymous):

3.316625

OpenStudy (aripotta):

that's not what i got

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

7.8?

OpenStudy (aripotta):

\[\sqrt{(2-(-4))^2 +(1-(-4))^2}=~ \sqrt{6^2 + 5^2}=~\sqrt{36+25}=~\sqrt{61}=~7.81\]

OpenStudy (aripotta):

yea

OpenStudy (anonymous):

ok thanks

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)